OpenStudy (dan815):

Let H be a hermitean(self-adjoint) operator on the 1-Qubit state space, show that e^(-iH) is unitary if a matrix is Unitary then U x U* = Identity If a H matrix is hermitean then H = H*

OpenStudy (dan815):

okay so im tempted to just say \[(e^{iH})^*=e^{-iH}\] dont know if that is right, but lets suppose that it is, then can i really just do \[e^{iH} ~x~e^{-iH} = e^{i(H-H)} = 1

OpenStudy (dan815):

\[e^{iH} ~*~e^{-iH} = e^{i(H-H)} = 1

OpenStudy (dan815):

where 1 is the identity mat

OpenStudy (empty):

I think you can say \[(e^{iH})^\dagger = e^{(iH)^\dagger} = e^{-i H}\] but I don't know if they expect you to go more in depth in this

OpenStudy (dan815):

kay cool lemme finish eating before the next qn bby

OpenStudy (empty):

actually even this might not be explicit enough, but it's better than what I wrote earlier. \[(e^{iH})^\dagger =\left( \sum_{n=0}^\infty \frac{(iH)^n}{n!}\right)^\dagger= \sum_{n=0}^\infty \frac{(-iH)^n}{n!}=e^{-iH}\]

OpenStudy (dan815):

ya

OpenStudy (dan815):

the explicit way i can think of needs an assuming that there is a P^T D P = H decomposition available

OpenStudy (dan815):

assumption*

OpenStudy (empty):

Nah, we're almost there let me just write it out and explain each step better

OpenStudy (dan815):

ok cool

OpenStudy (dan815):

also 1 more thing i guess it didnt matter for this question, but since H is a herm opp for a 1 qubbit space, does that mean H is a 2by 2 matrix

OpenStudy (dan815):

gotta show this to be true (A^n + B^n)*= A^n+B^n if A and B are hermetian

OpenStudy (dan815):

for this i can show it if A and B, are able to be written in P^T D P form but hmm

OpenStudy (dan815):

(symmetric)^n = still symmetric?

OpenStudy (empty):

Rewrite the e^x in taylor series: \[(e^{iH})^\dagger =\left( \sum_{n=0}^\infty \frac{(iH)^n}{n!}\right)^\dagger\] Transposing and complex conjugate both distribute over addition, so we can apply it to each term in the sum: \[\left( \sum_{n=0}^\infty \frac{(iH)^n}{n!}\right)^\dagger=\sum_{n=0}^\infty \left( \frac{(iH)^n}{n!}\right)^\dagger\] Do some algebra, I just rewrote H=IH and then multiplied the constants times the identity matrix on the left to separate them out. \[\sum_{n=0}^\infty \left( \frac{(iH)^n}{n!}\right)^\dagger = \sum_{n=0}^\infty \left[ \left( \frac{i^n}{n!}I \right) \left( H^n\right) \right]^\dagger\] Now when we transpose we have to write all the matrices in reverse order but since we have powers of the same matrix, each of them transposes simply: \[\sum_{n=0}^\infty \left[ \left( \frac{i^n}{n!}I \right) \left( H^n\right) \right]^\dagger = \sum_{n=0}^\infty\left( H^n\right)^\dagger \left( \frac{i^n}{n!}I \right) ^\dagger \] Now conjugate transpose them, but each of the \(H^\dagger = H\) and the only interesting thing about the conjugate transpose of a scalar multiplied by the identity is just \(i^*=-i\). \[\sum_{n=0}^\infty\left( H^n\right)^\dagger \left( \frac{i^n}{n!}I \right) ^\dagger = \sum_{n=0}^\infty H^n \frac{(-i)^n}{n!}I \] Scalars commute trivially with all matrices, and the identity matrix absorbs in so we just end up with: \[ \sum_{n=0}^\infty \frac{(-iH)^n}{n!}=e^{-iH}\] Done, no need to diagonalize anything, it's simple as that.

OpenStudy (dan815):

oh babyy i just realizedd that

OpenStudy (dan815):

(AB)^T = B^T * A^T

OpenStudy (dan815):

xD

OpenStudy (empty):

Don't worry about whether or not \((H^n)^\dagger = H^n\) Instead let's look at this: \[(ABC)^\dagger= C^\dagger B^\dagger A^\dagger\] now if we independently know \[C^\dagger = C\] and \[B^\dagger = B\] and \[A^\dagger = A\] then we can replace them all: \[(ABC)^\dagger = C^\dagger B^\dagger A^\dagger = CBA\] now if C=B=A=H now you have it \[(H^3)^\dagger = H^3\]

OpenStudy (dan815):

yepp

OpenStudy (empty):

ok cool lol I actually didn't know this until I typed all this out myself so it was a good exercise for me thanks lol

OpenStudy (dan815):

lol now i kinda realize why the T became dagger haha

OpenStudy (dan815):

they just added like 1 little line to T

OpenStudy (dan815):